Find 3 consecutive even integers such that of the sum of the smallest and the second is 5 more than 1/3 of the largest.

Answer by CubeyThePenguin(3113)

consecutive even integers: (x-2), x, (x+2)

(x-2) + x = (x+2)/3 + 5
2x - 2 = (x+2)/3 + 5
6x - 6 = x + 2 + 15
5x = 23
x = 23/5

The solution of x is not an integer, so this cannot happen.